# A Bimodal Metallicity Distribution Function in the Ultra-Faint Dwarf Galaxy Reticulum II

### v1

@@ 3.2.1 Zero-point calibration
It is not clear if the Ca II K calibration (Beers et al., 1999) is on the same scale as most of the literature on dwarf galaxies that uses the Calcium II triplet (CaT) calibration (Carrera et al., 2013). We compared our metallicities to medium-resolution spectroscopic metallicities with VLT/FLAMES from Simon et al. (2015) and Ji et al. (2023). Simon et al. (2015) used the CaT lines (8498, 8542, 8662Å) and Ji et al. (2023) used an Fe I line (6495Å) to measure . We have 9 stars cross-matched with Simon et al. (2015) and Ji et al. (2023), with 4 stars observed in both. The 13 metallicities used for our comparison are shown in Figure 6. There were 12 measurements more metal-poor and 1 measurement more metal-rich than in the literature, with a weighted average offset of dex. As literature values mostly use CaT spectroscopic metallicities, we apply a 0.29 dex shift to all of our reported values, including upper limits. Future work will focus on a full recalibration of this metallicity scale.
@@ 4.1 MDF of RetII
In Figure 7, we show the metallicity distribution function (MDF) of Reticulum II with 129 metallicities in the light gray filled histogram and the 38 upper limits stacked on top in the light gray line. Our MDF is compared to literature spectroscopic (Simon et al., 2015; Ji et al., 2023) and photometric (Fu et al., 2023) MDFs. We binned the histogram by the median metallicity uncertainty of 0.4 dex. The number of stars with spectroscopic metallicities in Ret II increased by 6.5x, making this the most populated spectroscopic or photometric UFD MDF currently in the literature. This was possible due to our spectroscopic metallicities of faint MSTO stars. By eye, we see a bimodality in the MDF that has not been seen before in any UFD, possibly because of the small number of stars observed. At , the uncertainties are lower by 0.2 dex, further motivating the investigation of bimodality, see Figure 8.
| Reference | ||||
@@
| Luna et al. 2025 | 167 | 38 | ||
| (This work) | ||||
@@ 4.2 Likelihood Analysis of MDF
We follow Kirby et al. (2011, 2013) for the likelihood functions of the detections and Ji et al. (2023) for the upper limits. We determined the best-fit model parameters, , by maximizing the likelihood function, , or the more computationally tractable optimization of the loglikelihood function, , where is the likelihood for one star, .
@@
We sampled the posterior with emcee (Foreman-Mackey et al., 2013) and initialized 64 walkers with 2000 burn-in iterations. We assumed log-flat priors on the parameters to limit the metallicity range explored. To determine a goodness of fit for each model, we use the maximum likelihood from the best-fit parameters and penalize any overfitting with extra parameters. We adopt the Akaike Information Criterion (AIC) (Akaike, 1974) described as:
@@ 4.3 Single Gaussian MDF
We determine the model parameters for a single Gaussian MDF, as previously done in the literature (e.g., Li et al. 2018). The Gaussian MDF has two parameters, mean metallicity, , and metallicity dispersion, :
@@
The MDF in Simon et al. (2015) was determined from 16 spectroscopic metallicities with the Calcium triplet calibration of Carrera et al. (2013). The MDF in Ji et al. (2023) is from 13 stars with an iron line fit using the radiative transfer and spectral synthesis code MOOG (Sneden, 1973). Fu et al. (2023) measured the metallicities for Ret II stars along the MSTO and main sequence (MS) using HST narrowband CaHK imaging and bolometric corrections from MIST. Their MDF has 76 photometric metallicities (including 26 upper limits), with a larger dispersion compared to spectroscopic studies. Their sample contains 1/3 of stars with , unlike spectroscopic MDFs, including ours. Table 2 shows the mean metallicities and dispersions of spectroscopic and photometric studies.
@@ 4.4 Analytical Chemical Evolution Models
Note. — Lower AIC means it is preferred over the Gaussian model. All models fit worse than a Gaussian model.
| Model | Uncertainty | AIC | ||||||
| Leaky Box + Gaussian | 5.1 | |||||||
| Gaussian + Gaussian | 0.76 | 3.5 | ||||||
| Leaky Box + Gaussian | 0.86 | 2.6 | ||||||
| Gaussian + Gaussian | 0.85 | 8.1 |
Note. — AIC is only comparable with the same data. AIC are relative to the single Gaussian model. A lower AIC indicates a better fit. The ∗ indicates a 90th percentile upper limit.
Fitting analytical models to the MDF allows us to learn about the gas flows in the system (Kirby et al., 2011). The Gaussian distribution parameters have no physical interpretation whereas chemical evolution (CE) model parameters do. Following Kirby et al. (2011), the different models we fit to the MDF are Leaky Box, Pre-Enriched, and Extra Gas models. The simplest model is the Leaky Box model, parameterized by the effective yield, . It describes a system that starts with a finite amount of pristine gas with , allows gas to leave, and does not accrete any other external gas (Pagel, 1997). A high effective yield could point to a system that retained more of its metals due to a deeper potential well or was more efficient at turning gas into stars. The Pre-Enriched model is the Leaky Box with the addition of an initial non-zero metallicity to the gas it starts with,. The Extra Gas model, also known as the best accretion model in Lynden-Bell (1975), is a system that both allows the gas to leave and accretes external metal-free gas. The amount of gas is described in the model with the parameter, M. The Extra Gas model becomes the Leaky Box model for M=1. Table 3 shows the best-fit parameters and the AIC value, which is compared to the AIC value of a single Gaussian MDF (AIC ). The CE models are not preferred over a single Gaussian model. Figure 9 shows the CE models compared to the Gaussian MDF.
@@ 4.5 Bimodal Model Fitting
The uncertainties in the Beers et al. (1999) calibration are not well characterized, affecting of our sample that are dominated by systematic uncertainties. It is unclear whether the sources of the uncertainty are from a zero-point calibration, stellar parameter trends, or intrinsic variability of the metallicities. An updated calibration could potentially reduce the systematic uncertainty. To explore a best-case scenario, we run each model separately with the total metallicity uncertainty ( 0.4) and with only the spectroscopic metallicity uncertainty ( 0.2). Appendix B shows the posterior distributions of the model parameters and Table 4 shows the best-fit parameter values for each mixture model and their AIC values. The AIC values are relative to the Gaussian MDF (AIC ). Figure 10 shows the mixture model fits convolved with the typical uncertainty of dex. The AIC values indicate a Leaky Box + Gaussian MM is only preferred over a single Gaussian model when using the spectroscopic uncertainties. The Gaussian + Gaussian MM is preferred in both uncertainty cases. It is slightly preferred with the total uncertainty, but more strongly preferred with the spectroscopic uncertainty, supporting the presence of bimodality in the MDF.
We perform a Kolmogorov–Smirnov (KS) test on the MS, MSTO, and sub-RGB/RGB regions to test our confidence in the bimodality by comparing the cumulative metallicity distributions of our entire sample with those of the different regions. Shown in Figure 11, the p-values suggest that the entire sample and the MS, MSTO, and sub-RGB/RGB regions are each derived from the same parent distribution. The rightmost plots show the CDF and CMD including stars from Simon et al. (2015), increasing our confidence in the bimodality of the sub-RGB/RGB region from the high p-value.
In Figure 12, we show the spatial distribution of our sample colored by metallicity. We plot the CDFs of metal-poor stars ([Fe/H] ) and of the metal-rich stars ([Fe/H] ). We find no spatial separation between the metal-poor and metal-rich populations.
@@ 5.1 Interpreting the Bimodality of the Ret II MDF
We report the first bimodal MDF seen in any UFD 111The Segue 1 UFD was reported to have a wide dispersion or multi-modal metallicity distribution from the range of 7 RGB metallicities spanning more than 2 dex (Frebel et al., 2014; Webster et al., 2016). Bissonette et al. (in prep.) has recently confirmed that Segue 1 has a Gaussian MDF with .. There is a peak at with of the stars and another peak at with of the stars. The MDF has a metallicity gap of dex. Ting & Ji (2024) proposed a method to predict the observed metallicity gap in an MDF, due to quiescent periods between star formation events, for a range of dwarf galaxy masses. The stars from the first burst undergo Type Ia supernovae and eject iron, enriching the gas and resulting in the subsequent burst of stars forming with a higher metallicity. Using the SFH in Simon et al. (2023), they determined a 1.02 dex for Ret II, consistent with the gap in our observed MDF, dex. This supports their theory of predicting metallicity gaps in dwarf galaxies.
@@
The current SFH for Reticulum II in Simon et al. (2023) prefers a two-burst star formation model. The best-fit instantaneous burst model has a burst at 14.1 Gyr containing 87.5% of stars and a second burst 3.4 Gyr later forming the final 12.5% of stars. They constrain SFH models by selecting isochrones that follow the single-mode MDF in Simon et al. (2015). The percentage of stars in each burst from the SFH is consistent with the results of our bimodal MDF analysis, where we have a mixing fraction of when using total uncertainties, and with only spectroscopic uncertainties we have . Further constraining the models going into the SFH to follow a bimodal MDF would result in a more accurate distribution of isochrones, and potentially the SFH.
@@
As an initial estimate, we determined the age-metallicity relation (AZR) using the cumulative SFH in Simon et al. (2023) interpolated with 0.1 steps to the CDF of our bimodal metallicity distribution, shown in Figure 13. The AZR combines our MDF with the SFH by assuming that metallicity increases monotonically with time, with the metal-rich peak consisting of younger stars. The AZR suggests that the metallicity gap of 1 dex is related to an age gap of 3 Gyr, and therefore that there is star formation after reionization.
@@
Other independent analyses of the formation history of Reticulum II indicate similar results. Alexander et al. (2023) predicted a bimodal metallicity distribution of Ret II from modeling inhomogeneous chemical evolution. They interpret the higher metallicity mode as a consequence of implementing a delay-time distribution for Type Ia SNe in their modeling. Ji et al. (2023) also found independent evidence of bursty star formation in Reticulum II from well-mixed Barium abundances in r-process enriched stars.
@@
Although we prefer the two-burst star formation scenario, the accretion of multiple progenitor halos could also result in a bimodal MDF. Simulations suggest that UFDs may form their stars in multiple smaller halos from different environments that eventually merge (Simpson et al., 2013; Jeon et al., 2017) and can be dominated by late-time dry mergers (Rey et al., 2019; Andersson et al., 2025). The progenitors evolve to have distinct MDFs that are imprinted on the overall MDF once accreted (Ko et al., 2024). For example, a halo at can have accreted many progenitor halos with varying metallicities. This is a possible interpretation for the bimodal MDF. Although Ko et al. (2024) did not find a bimodality in any resulting halo MDF, we note that it could be possible if there were not many accretions that could wash out the bimodality in the final MDF. The accreted population is more likely to be more metal-poor, however, to agree with the Ret II MDF, the accreted population comprising of the minority of stars is more metal-rich. This scenario could be distinguishable from the two-burst star formation scenario from the AZR and the spatial distribution. Figure 14 shows the expected results of the spatial metallicity distribution and AZR for the two scenarios. We investigated the spatial distribution of our sample in Figure 12 and found no spatial separation between the metal-rich and metal-poor populations suggesting a preference for the two-burst scenario, but a full AZR with independent measurements of age and metallicity would be decisive.
@@ 5.2 Implications of the Formation History and Evolution of Ret II and other UFDs
Ret II was discovered to have experienced a rare and prolific r-process event that produced large amounts of r-process elements such as Europium and Barium (Ji et al., 2016a, b; Roederer et al., 2016). There is a current debate about the possible sites for r-process nucleosynthesis (Frebel & Ji, 2023). The two most likely astrophysical sites are neutron star mergers (NSM) and rare core-collapse supernovae (CCSNe), which occur at drastically different delay times, with the former taking up to billions of years and the latter all occurring after a few million years. Simon et al. (2023) constrained the timing of the nucleosynthetic event by studying the extended star formation history using that it must have occurred after 28 of non-r-process-enhanced stars had formed (Ji et al., 2023) and found that the event occurred within 500 200 Myr of the formation of the first stars in Ret II. The AZR in Figure 13 shows that the second burst of stars must all be r-process enriched. With the second burst of stars having higher [Fe/H], we should see a decrease in [r/Fe] when there is an increase in Fe. A possible indication of this trend in the literature is that the star DES J033548540349 was relatively metal-rich with but somewhat less enhanced in neutron-capture abundances than the metal-poor stars in Ji et al. (2016b).
Star formation on the smallest scales is quenched or suppressed during the epoch of reionization (Bullock et al., 2000; Brown et al., 2014; Weisz et al., 2014; Jeon et al., 2017; Rodriguez Wimberly et al., 2019), however, it is unclear if this effect is homogeneous for all UFDs. The SFHs of Magellanic Cloud (MC) satellites are suggested to continue 600 Myr longer than the Milky Way (MW) satellites, due to a weaker local ionization field during reionization, emphasizing the importance of environmental effects on star formation in UFDs (Sacchi et al., 2021; Kim et al., 2023; Durbin et al., 2025). Savino et al. (2023) shows that compared to MW satellites, M31 satellites continue to form stars longer after reionization, concluding that M31 UFDs are not as affected by reionization. They note that the M31 UFDs in their study are more massive than the MW satellites, which could affect the amount of gas retained. A study of Pegasus W, an isolated UFD outside the virial radius of M31, also shows the impact of environment on star formation as it formed of its stars after (McQuinn et al., 2023). These studies suggest that reionization is not homogeneously quenching star formation, and that environment is also an important factor. In this study, we find that reionization did not completely quench the star formation in Reticulum II, with a stellar mass of , which simulations predict is below the minimum mass threshold of galaxies massive enough to retain gas and form stars after reionization (e.g., Kravtsov et al., 2004). The AZR in Figure 13 shows a second burst post-reionization, suggesting the reignition of star formation. Therefore, Ret II may have finally self-quenched from internal mechanisms such as stellar or supernova feedback, not by reionization (e.g., Gallart et al., 2021). With more well-populated MDFs in UFDs, we can detect metallicity gaps related to gaps in age, revealing the importance of reionization versus internal quenching mechanisms.

### v2

@@ 3.2.1 Zero-point calibration
It is not clear if the Ca II K calibration (Beers+1999) is on the same scale as most of the literature on dwarf galaxies that uses the Calcium II triplet (CaT) calibration (Carrera2013). For our sample, the CaT calibration is not ideal due to weaker lines in hotter MSTO stars. Additionally, there is no published CaT calibration for MS stars. We compared our metallicities to medium-resolution spectroscopic metallicities with VLT/FLAMES from Simon+2015 and Ji+2023. Simon+2015 used the CaT lines (8498, 8542, 8662Å) and Ji+2023 used an Fe I line (6495Å) to measure . We have 9 stars cross-matched with Simon+2015 and Ji+2023, with 4 stars observed in both. The 13 metallicities used for our comparison are shown in Figure 6. We note that Ji+2023 uses Dartmouth isochrones (Dotter2008), so metallicities may not be directly comparable. There were 12 measurements more metal-poor and 1 measurement more metal-rich than in the literature, with a weighted average offset of dex. As literature values mostly use CaT spectroscopic metallicities, we apply a 0.29 dex shift to all of our reported values, including upper limits. Future work will focus on a full recalibration of this metallicity scale.
@@ 4.1 MDF of RetII
In Figure 7, we show the metallicity distribution function (MDF) of Reticulum II with 129 metallicities in the light gray filled histogram and the 38 upper limits stacked on top in the light gray line. Our MDF is compared to literature spectroscopic (Simon+2015; Ji+2023) and photometric (Fu+2023) MDFs. We binned the histogram by the median metallicity uncertainty of 0.4 dex. The number of stars with spectroscopic metallicities in Ret II increased by 6.5x, making this the most populated spectroscopic or photometric UFD MDF currently in the literature. This was possible due to our spectroscopic metallicities of faint MSTO stars. By eye, we see a bimodality in the MDF that has not been seen before in any UFD, possibly because of the small number of stars observed. At , the uncertainties are lower by 0.2 dex, further motivating the investigation of bimodality, see Figure 8.
[…]
@@
| Luna et al. 2025 | ||||
| This work, All | 167 | 38 | ||
| This work, Det. only | 129 | 0 | ||
@@ 4.2 Likelihood Analysis of MDF
We follow Kirby+2011; Kirby2013 for the likelihood functions of the detections and Ji+2023 for the upper limits. We determined the best-fit model parameters, , by maximizing the likelihood function, , or the more computationally tractable optimization of the loglikelihood function, , where is the likelihood for one star, .
@@
We sampled the posterior with emcee (emcee2013) and initialized 64 walkers with 2000 burn-in iterations. We assumed log-flat priors on the parameters to limit the metallicity range explored. To determine a goodness of fit for each model, we use the maximum likelihood from the best-fit parameters and penalize any overfitting with extra parameters. We adopt the Akaike Information Criterion (AIC) (Akaike1974) described as:
@@ 4.3 Single Gaussian MDF
We determine the model parameters for a single Gaussian MDF, as previously done in the literature (e.g., Li2018). The Gaussian MDF has two parameters, mean metallicity, , and metallicity dispersion, :
@@
Table 1 shows the mean metallicities and dispersions of spectroscopic and photometric studies. The MDF in Simon+2015 was determined from 16 spectroscopic metallicities with the Calcium triplet calibration of Carrera2013. The MDF in Ji+2023 is from 13 stars with an iron line fit using the radiative transfer and spectral synthesis code MOOG (Sneden1973). Fu+2023 measured the metallicities for Ret II stars along the MSTO and main sequence (MS) using HST narrowband CaHK imaging and bolometric corrections from MIST. Their MDF has 76 photometric metallicities (including 26 upper limits), with a larger dispersion compared to spectroscopic studies. Their sample contains 1/3 of stars with , unlike spectroscopic MDFs, including ours.
@@ 4.4 Analytical Chemical Evolution Models
Lower AIC means it is preferred over the Gaussian model. All models fit worse than a Gaussian model.
[…]
AIC is only comparable with the same data. AIC are relative to the single Gaussian model. A lower AIC indicates a better fit. The ∗ indicates a 90th percentile upper limit.
Fitting analytical models to the MDF allows us to learn about the gas flows in the system (Kirby+2011). The Gaussian distribution parameters have no physical interpretation whereas chemical evolution (CE) model parameters do. Following Kirby+2011, the different models we fit to the MDF are Leaky Box, Pre-Enriched, and Extra Gas models. The simplest model is the Leaky Box model, parameterized by the effective yield, . It describes a system that starts with a finite amount of pristine gas with , allows gas to leave, and does not accrete any other external gas (Pagel1997). A high effective yield could point to a system that retained more of its metals due to a deeper potential well or was more efficient at turning gas into stars. The Pre-Enriched model is the Leaky Box with the addition of an initial non-zero metallicity to the gas it starts with,. The Extra Gas model, also known as the best accretion model in Lynden-Bell1975, is a system that both allows the gas to leave and accretes external metal-free gas. The amount of gas is described in the model with the parameter, M. The Extra Gas model becomes the Leaky Box model for M=1. Table 2 shows the best-fit parameters and the AIC value, which is compared to the AIC value of a single Gaussian MDF (AIC ). The CE models are not preferred over a single Gaussian model. Figure 9 shows the CE models compared to the Gaussian MDF.
@@ 4.5 Bimodal Model Fitting
The uncertainties in the Beers+1999 calibration are not well characterized, affecting of our sample that are dominated by systematic uncertainties. It is unclear whether the sources of the uncertainty are from a zero-point calibration, stellar parameter trends, or intrinsic variability of the metallicities. An updated calibration could potentially reduce the systematic uncertainty. To explore a best-case scenario, we run each model separately with the total metallicity uncertainty ( 0.4) and with only the spectroscopic metallicity uncertainty ( 0.2). Appendix B shows the posterior distributions of the model parameters and Table 3 shows the best-fit parameter values for each mixture model and their AIC values. The AIC values are relative to the Gaussian MDF (AIC ). Figure 10 shows the mixture model fits convolved with the typical uncertainty of dex. The AIC values indicate a Leaky Box + Gaussian MM is only preferred over a single Gaussian model when using the spectroscopic uncertainties. The Gaussian + Gaussian MM is preferred in both uncertainty cases. It is slightly preferred with the total uncertainty, but more strongly preferred with the spectroscopic uncertainty, supporting the presence of bimodality in the MDF.
We perform a Kolmogorov–Smirnov (KS) test on the MS, MSTO, and sub-RGB/RGB regions to test our confidence in the bimodality by comparing the cumulative metallicity distributions of our entire sample with those of the different regions. Shown in Figure 11, the p-values suggest that the entire sample and the MS, MSTO, and sub-RGB/RGB regions are each derived from the same parent distribution. The rightmost plots show the CDF and CMD including stars from Simon+2015, increasing our confidence in the bimodality of the sub-RGB/RGB region from the high p-value.
Mergers could affect the spatial separation of metal-poor and metal-rich populations, as possibly seen in the Sculptor dwarf galaxy (Benitez-Llambay2016; Arroyo-Polonio2024; Barbosa2025). In Figure 12, we show the spatial distribution of our sample colored by metallicity. We plot the CDFs of metal-poor stars ([Fe/H] ) and of metal-rich stars ([Fe/H] ). Visually, it seems like the most metal-poor stars could be biased to the northern half of Ret II. However, we find no significant spatial separation in the metal-poor and metal-rich populations, along either the radial or the directions.
@@ 5.1 Interpreting the Bimodality of the Ret II MDF
We report the first bimodal MDF seen in any UFD 111The Segue 1 UFD was reported to have a wide dispersion or multi-modal metallicity distribution from the range of 7 RGB metallicities spanning more than 2 dex (Frebel2014; Webster2016). Bissonette et al. (in prep.) has recently confirmed that Segue 1 has a Gaussian MDF with .. There is a peak at with of the stars and another peak at with of the stars. The MDF has a metallicity gap of dex. Ting&Ji2024 proposed a method to predict the observed metallicity gap in an MDF, due to quiescent periods between star formation events, for a range of dwarf galaxy masses. The stars from the first burst undergo Type Ia supernovae and eject iron, enriching the gas and resulting in the subsequent burst of stars forming with a higher metallicity. Using the SFH in Simon+2023, they determined a 1.09 dex for Ret II, consistent with the gap in our observed MDF, dex. This supports their theory of predicting metallicity gaps in dwarf galaxies.
@@
The current SFH for Reticulum II in Simon+2023 prefers a two-burst star formation model. The best-fit instantaneous burst model has a burst at 14.1 Gyr containing 87.5% of stars and a second burst 3.4 Gyr later forming the final 12.5% of stars. They constrain SFH models by selecting isochrones that follow the single-mode MDF in Simon+2015. The percentage of stars in each burst from the SFH is consistent with the results of our bimodal MDF analysis, where we have a mixing fraction of when using total uncertainties, and with only spectroscopic uncertainties we have . Further constraining the models going into the SFH to follow a bimodal MDF would result in a more accurate distribution of isochrones, and potentially the SFH.
@@
As an initial estimate, we determined the age-metallicity relation (AZR) using the cumulative SFH in Simon+2023 interpolated with 0.1 steps to the CDF of our bimodal metallicity distribution, shown in Figure 13. The AZR combines our MDF with the SFH by assuming that metallicity increases monotonically with time, with the metal-rich peak consisting of younger stars. The AZR suggests that the metallicity gap of 1 dex is related to an age gap of 3 Gyr, and therefore that there is star formation after reionization.
@@
Other independent analyses of the formation history of Reticulum II indicate similar results. Alexander+2023 predicted a bimodal metallicity distribution of Ret II from modeling inhomogeneous chemical evolution. They interpret the higher metallicity mode as a consequence of implementing a delay-time distribution for Type Ia SNe in their modeling. Ji+2023 also found independent evidence of bursty star formation in Reticulum II from well-mixed Barium abundances in r-process enriched stars.
@@
Although we prefer the two-burst star formation scenario, the accretion of multiple progenitor halos could also result in a bimodal MDF. Simulations suggest that UFDs may form their stars in multiple smaller halos from different environments that eventually merge (Simpson2013; Jeon2017) and can be dominated by late-time dry mergers (Rey2019; Andersson2025). The progenitors evolve to have distinct MDFs that are imprinted on the overall MDF once accreted (Ko2024). For example, a halo at can have accreted many progenitor halos with varying metallicities. This is a possible interpretation for the bimodal MDF. Although Ko2024 did not find a bimodality in any resulting halo MDF, we note that it could be possible if there were not many accretions that could wash out the bimodality in the final MDF. The accreted population is more likely to be more metal-poor, however, to agree with the Ret II MDF, the accreted population comprising of the minority of stars is more metal-rich. This scenario could be distinguishable from the two-burst star formation scenario from the AZR and the spatial distribution. Figure 14 shows the expected results of the spatial metallicity distribution and AZR for the two scenarios. We investigated the spatial distribution of our sample in Figure 12 and found no spatial separation between the metal-rich and metal-poor populations suggesting a preference for the two-burst scenario, but a full AZR with independent measurements of age and metallicity would be decisive.
Another possible scenario that could affect the MDF is related to gas accretion of the intergalactic medium (IGM). Ahvazi2024 introduced an IGM metallicity model in their simulations, reproducing the observed mass-metallicity relation (MZR) at small scales. However, Wheeler2025 found that the amount of IGM accretion is not sufficient to affect the MDF. While more work is needed to study UFD MDFs, we expect the AZR to be distinguishable from the other scenarios.
5.2 r-process in Ret II
Ret II was discovered to have experienced a rare and prolific r-process event that produced large amounts of r-process elements such as Europium and Barium (Ji+2016a; Ji+2016c; Roederer2016). There is a current debate about the possible sites for r-process nucleosynthesis (Frebel&Ji2023). The two most likely astrophysical sites are neutron star mergers (NSM) and rare core-collapse supernovae (CCSNe), which occur at drastically different delay times, with the former taking up to billions of years and the latter all occurring after a few million years. Simon+2023 constrained the timing of the nucleosynthetic event by studying the extended star formation history using that it must have occurred after 28 of non-r-process-enhanced stars had formed (Ji+2023) and found that the event occurred within 500 200 Myr of the formation of the first stars in Ret II. The AZR in Figure 13 shows that the second burst of stars must all be r-process enriched. With the second burst of stars having higher [Fe/H], we should see a decrease in [r/Fe] when there is an increase in Fe. A possible indication of this trend in the literature is that the star DES J033548540349 was relatively metal-rich with but somewhat less enhanced in neutron-capture abundances than the metal-poor stars in Ji+2016c.
@@ 5.3 Implications of the Formation History and Evolution of Ret II and other UFDs
Star formation on the smallest scales is quenched or suppressed during the epoch of reionization (Bullock2000; Brown2014; Weisz2014; Jeon2017; RodriguezWimberly2019), however, it is unclear if this effect is homogeneous for all UFDs. The SFHs of Magellanic Cloud (MC) satellites are suggested to continue 600 Myr longer than the Milky Way (MW) satellites, due to a weaker local ionization field during reionization, emphasizing the importance of environmental effects on star formation in UFDs (Sacchi+2021; Kim2023; Durbin2025). Savino2023 shows that compared to MW satellites, M31 satellites continue to form stars longer after reionization, concluding that M31 UFDs are not as affected by reionization. They note that the M31 UFDs in their study are more massive than the MW satellites, which could affect the amount of gas retained. A study of Pegasus W, an isolated UFD outside the virial radius of M31, also shows the impact of environment on star formation as it formed of its stars after (McQuinn2023_PegasusW). These studies suggest that reionization is not homogeneously quenching star formation, and that environment is also an important factor. In this study, we find that reionization did not completely quench the star formation in Reticulum II, with a stellar mass of , which simulations predict is below the minimum mass threshold of galaxies massive enough to retain gas and form stars after reionization (e.g., Kravtsov2004). The AZR in Figure 13 shows a second burst post-reionization, suggesting the reignition of star formation. Therefore, Ret II may have finally self-quenched from internal mechanisms such as stellar or supernova feedback, not by reionization (e.g., Gallart2021). With more well-populated MDFs in UFDs, we can detect metallicity gaps related to gaps in age, revealing the importance of reionization versus internal quenching mechanisms.